The length of one of the equal legs of an isosceles triangle is 8 cm less than 4 times the length of the base. If the perimeter is 29 cm, find the length of one of the equal legs.
Answer by ptaylor(1894)


Let x=length of base
then 4x-8=length of one of the equal legs

Now we are told that:
x+2(4x-8)=29 get rid of parens
x+8x-16=29 add 16 to both sides
x+8x-16+16=29+16 collect like terms
9x=45 divide both sides by 9
x=5 cm---------------------length of base
4x-8=4(5)-8=12cm------------length of one of the equal legs
CK
12+12+5=29
29=29
Hope this helps---ptaylor